A pigeon can fly at a speed of 84 kilometers per hour. How long does it take the pigeon to fly 7 kilometers?
Found 2 solutions by rfer, iamverybadinmath:
Answer by rfer(16322)

84/1=7/x
84x=7
x=7/84
x=5 minutes

Answer by iamverybadinmath(9)

7 kilometers/84kilometers per hour
=1/5
=0.2hour
convert hour to minutes
=0.2x60
=12 minutes